OpenStudy (anonymous):

Trevor's dad decided to get him a cell phone for his birthday. Cell Plus has a plan that costs $24.00 per month plus an additional $0.06 per minute. Cell Best has a plan that costs $30.00 per month plus an additional $0.05 per minute. How many minutes can Trevor talk and both cell plans cost the same amount? 600 60 300 120

OpenStudy (anonymous):

600 minutes Convert $ to pennies and solve the following for x: 2400+6m=3000+5 m m = 600